OpenStudy (anonymous):

Find dy/dx if f(x) = (x+1)^(2x)

OpenStudy (solomonzelman):

Logarithmic differentiation

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y=(x+1)^{2x} }\) hit both sides with a natural log, and then differentiate on both sides.

OpenStudy (solomonzelman):

Or,

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y=(x+1)^{2x} =e^{\ln(x+1)^{2x}}=e^{2x\ln(x+1)}}\)

OpenStudy (owlcoffee):

Here we can make use of the derivation rule: \[f(x)=A^n \iff f'(x)=(n)A ^{n-1}\] Now, this will be okay, but thing is, what we have in front of us is what we call a composite function, since the racial is also a function, we might want to conveniently transport it to the numerator by applying a logarithm. Or, as I would do it: \[f(x) = (x+1)^{2x} \iff f(x)=e ^{\ln(x+1)^{2x}}\] Which gets traduced to: \[f(x)=e ^{(2x)\ln(x+1) }\] And that is most easy to derive, after all, the drivative of: \[h(x)=e ^{f(x)} \iff h'(x)= e ^{f(x)}.f'(x)\]

OpenStudy (solomonzelman):

let \(\color{#000000 }{ \displaystyle f }\) and \(\color{#000000 }{ \displaystyle g }\) be differentiable functions of x, and let \(\color{#000000 }{ \displaystyle z=f^g }\). We can do a little algebra; \(\color{#000000 }{ \displaystyle z=f^g}\) \(\color{#000000 }{ \displaystyle z=e^{\large \ln(f^g)} }\) \(\color{#000000 }{ \displaystyle z=e^{\large ~g\ln(f)} }\) Then the derivative \(\color{#000000 }{ \displaystyle \frac{dz}{dx}=e^{\large ~g\ln(f)}\times \left[ g'\ln(f)+g\frac{f'}{f}\right] }\)

OpenStudy (solomonzelman):

In your case, as you will see (when you come online). f = (x+1) g = 2x

OpenStudy (solomonzelman):

saying over the same thing the third time in a different form.

OpenStudy (trojanpoem):

Or use the damn formula for function power function: y = n^y y' = y n^(y-1) * n' + n^y ln(n) * y' the same as taking ln, differentiating and then multiplying times the y ( as Solomon said).